A shunt motor armature resistance 0.08Ω is connected to 120V mains. With 50A in the armature, what are the back emf and the mechanical power developed within the armature?

1
Expert's answer
2021-12-13T10:34:38-0500

"r=\frac UI=2.4~\Omega,"

"\epsilon=I(R+r)=2.62~kV,"

"P=\varepsilon I=131~kW."